# Efficient Design for Multi-user Downlink Beamforming with Reconfigurable Intelligent Surface

### v2

@@ VII Simulation Results
We consider a downlink RIS-assisted transmission scenario with a BS and an RIS. The locations of the BS and the RIS using the $(x,y,z)$-coordinates in meters are $(0,0,0)$ and $(70,70,0)$, respectively. The RIS is a rectangular surface with the array elements placed on the $(y,z)$-plane. The users are located randomly within a circle centered at the RIS with a radius of $20$ m on the $(x,y)$-plane. We set the target SINR $\gamma_{ik}=\gamma$, $\forall i,k$. Unless specified otherwise, our default system setup is given as $N=16$ antennas at the BS, and $M=10\times 10$ reflective elements at the RIS, $G=2$ groups, $K_{i}=K=2$ users, $\forall i$.
@@
We assume a line-of-sight (LOS) channel between the BS and the RIS and model ${\bf H}_{\text{r}}$ as a Rician fading channel matrix given by ${\bf H}_{\text{r}}=\beta^{\text{BR}}\big{(}\sqrt{\frac{K_{\text{r}}}{1+K_{\text{r}}}}{\bf H}_{\text{r}}^{\text{LOS}}+\sqrt{\frac{1}{1+K_{\text{r}}}}{\bf H}_{\text{r}}^{\text{NLOS}}\big{)}$, where Rician factor $K_{\text{r}}=10$, $\beta^{\text{BR}}$ is the path gain modeled as $\beta^{\text{BR}}\text{[dB]}=-30-22\log 10(d^{\text{B-R}})$ with $d^{\text{B-R}}$ being the BS-RIS distance in meters, ${\bf H}_{\text{r}}^{\text{NLOS}}$ is the non-line-of-sight (NLOS) component modeled as $[{\bf H}_{\text{r}}^{\text{NLOS}}]_{ij}\sim{\cal C}{\cal N}(0,{\bf I})$, and ${\bf H}_{\text{r}}^{\text{LOS}}$ is the LOS component. The LOS component ${\bf H}_{\text{r}}^{\text{LOS}}$ is a function of the BS and RIS locations and is modeled as
${\bf H}_{\text{r}}^{\text{LOS}}={\bf b}_{\text{RIS}}(\psi_{2},\theta_{2}){\bf b}_{\text{BS}}(\psi_{1},\theta_{1})^{H}$ (43)
@@
where $\psi_{1}$ and $\theta_{1}$ are the azimuth and elevation angles of departure (AoD) from the BS to the IRS, and ${\bf b}_{\text{BS}}(\psi_{1},\theta_{1})$ is the BS antenna array steering vector, with its $n$th entry as
$[{\bf b}_{\text{BS}}(\psi_{1},\theta_{1})]_{n}=e^{\frac{2\pi(n-1)\Delta_{\text{BS}}}{\lambda_{c}}}\cos(\psi_{1})\cos(\theta_{1}),\ n=1,\ldots,N$
@@
where $\Delta_{\text{BS}}$ is the distance between two adjacent BS antennas, $\lambda_{c}$ is the carrier wavelength, and $\cos(\psi_{1})\cos(\theta_{1})=\frac{x_{\text{RIS}}-x_{\text{BS}}}{d^{\text{B-R}}}$. We assume $\Delta_{\text{BS}}=\frac{\lambda_{c}}{2}$. Similarly, $\psi_{2}$ and $\theta_{2}$ are the azimuth and elevation angles of arrival (AoA) from the BS to the RIS, and ${\bf b}_{\text{RIS}}(\psi_{2},\theta_{2})$ is the RIS element streaming vector, with the $m$th entry, for $m\in{\cal M}$, given by
$[{\bf b}_{\text{RIS}}(\psi_{2},\theta_{2})]_{m}=e^{\frac{2\pi\Delta_{\text{RIS}}}{\lambda_{c}}}[y_{m}\sin(\psi_{2})\cos(\theta_{2})+z_{m}\sin(\theta_{2})]$
@@
where $\Delta_{\text{RIS}}$ is the distance between two adjacent RIS elements, $y_{m}=\text{mod}(m-1,M_{y})$ with $M_{y}$ being the number of RIS elements along the $y$-axis, $z_{m}=\lfloor m/M_{y}\rfloor+1$, $\sin(\psi_{2})\cos(\theta_{2})=\frac{y_{\text{BS}}-y_{\text{RIS}}}{d_{\text{B-R}}}$, and $\sin(\theta_{2})=\frac{z_{\text{BS}}-z_{\text{RIS}}}{d_{\text{B-R}}}$. We assume $\Delta_{\text{RIS}}=\frac{\lambda_{c}}{2}$.
@@
We consider a NLOS channel between the RIS and each user modeled as Rayleigh fading. The channel ${\bf h}^{\text{r}}_{ik}$ between the RIS and user $k$ in group $i$ is generated i.i.d as ${\bf h}^{\text{r}}_{ik}\sim{\cal C}{\cal N}({\bf 0},\beta^{\text{r}}_{ik}{\bf I})$, where $\beta^{\text{r}}_{ik}$ is the path gain modeled as $\beta^{\text{r}}_{ik}[\text{dB}]=-30-22\log 10(d^{\text{R-U}}_{ik})$, with $d^{\text{R-U}}_{ik}$ being the distance between the RIS and the user. We also model the channel between BS and user $k$ in group $i$ as Rayleigh fading, which is generated i.i.d as ${\bf h}^{\text{d}}_{ik}\sim{\cal C}{\cal N}({\bf 0},\beta^{\text{d}}_{ik}{\bf I})$, where $\beta^{\text{d}}_{ik}$ is the path gain modeled as $\beta^{\text{d}}_{ik}[\text{dB}]=-32.6-36.7\log 10(d^{\text{B-U}}_{ik})$, with $d^{\text{B-U}}_{ik}$ being the respective BS-user distances. The receiver noise power is set to $\sigma^{2}=-100~{}\text{dBm}$.
@@
VII-A Convergence Behavior of the Proposed Algorithms
We first study the convergence behavior of our proposed AMBF algorithm in Algorithm 1 for the QoS problem ${\cal P}_{o}$, which alternatingly solves two multicast subproblems at the BS and the RIS using our proposed fast algorithms. Fig. 0(a) shows the convergence behavior of solving subproblem $\widetilde{{\cal S}}_{{\bf w}}$ using PSA in a specific iteration of the AMBF algorithm. The convergence behavior for three random channel realizations are shown. We see PSA converges after $1000\sim 2000$ iterations. Note that the closed-form updates for PSA in each iteration is computationally cheap, and thus, the overall computation time for $1000\sim 2000$ iterations is fast. Furthermore, denote the total transmit power at the BS as $P_{\text{tot}}=\sum_{i=1}^{G}\|{\bf w}_{i}\|^{2}$. Fig. 0(b) shows the trajectory of $P_{\text{tot}}$ over iterations by Algorithm 1 for three random channel realizations. We observe that our proposed AMBF approach converges within about $6\sim 12$ iterations, which is relatively fast, especially given that we use fast algorithms for computing the solution to each subproblem.
(a) The subproblem $\widetilde{{\cal S}}_{{\bf w}}$.
(b) The overall AMBF Algorithm.
Figure 1: Convergence behavior of the AMBF algorithm in Algorithm 1 ($N=16,M=100$, $(G,K)=(2,2)$, $\gamma=10$ dB).
Fig. 2 shows the convergence behavior of the MMF problem ${\cal S}_{o}$ solved by Algorithm 2 under three random channel realizations. We see that for two channel realizations, Algorithm 2 converges very fast within just a few iterations. Even for a worse scenario under the third channel realization, Algorithm 2 converges after $800$ iterations. Note that the PSA updates in each iteration is closed-form and computationally cheap.
Figure 2: Convergence behavior of the Algorithm 2 ($N=30$, $M=100$, $(G,K)=(2,2)$, $\gamma=10$ dB, $P=10$ dB).
VII-B Performance Comparison for the QoS Problem
We now evaluate the performance of our proposed AMBF algorithm in Algorithm 1. For performance comparison, we consider the following methods:
@@ VII-B Performance Comparison for the QoS Problem
    Direct AO [19]: Apply the AO approach directly to the QoS problem ${\cal P}_{o}$, where the subproblem w.r.t. ${\bf e}$ is a feasibility problem with the SINR constraints in (3b) and the unit-modulus constraints in (3c).
  4. iv)
    AMBF-SDR: Apply Algorithm 1, except that instead of solving subproblem $\tilde{{\cal S}}_{\bf w}$ w.r.t. ${\bf e}$, we relax ${\cal S}_{{\bf w}}^{\text{eq}}$ by dropping the constraint in (10c) and solve the relaxed problem:
    ${\cal S}_{{\bf w}}^{\text{Relaxed}}:\max_{{\bf e}}\min_{i,k}$ $\ \frac{\textrm{SINR}_{ik}}{\gamma_{ik}}$ s.t. $|e_{m}|^{2}\leq 1,m\in{\cal M},$
    We solve ${\cal S}_{{\bf w}}^{\text{Relaxed}}$ by using its epigraph form and applying SDR with a bi-section search for the objective value (i.e., let ${\bf X}={\bf e}{\bf e}^{H}$ and drop the rank-1 condition). Then, we recover ${\bf e}$ from ${\bf X}$ using the Gaussian randomization method.
@@
    AMBF-SCA: Similar to AMBF-SDR that apply Algorithm 1, except that we solve ${\cal S}_{{\bf w}}^{\text{Relaxed}}$ by SCA with bi-section search over $t$. In particular, we convexify ${\cal S}_{{\bf w}}^{\text{Relaxed}}$ into a sequence of convex approximation problems and solved them using a standard convex solver.
  6. vi)
    IP Method: Apply MATLAB non-linear optimization solver Fmincon to ${\cal P}_{o}$. It uses the interior-point (IP) algorithm to solve ${\cal P}_{o}$.
Figure 3: Total transmit power $P_{\text{tot}}$ vs. $M$ ($N=16$, $(G,K)=(2,2)$, $\gamma=10$ dB).
@@
$M$
25
50
100
200
300
RIS-SDR
24.73
46.95
84.68
672
1743
RIS-SCA
23.12
38.52
62.17
436
1349
IP Method
3.25
10.73
58.33
796
2062
Algorithm 1
2.19
3.12
5.22
8.83
13.62
VII-B1 Performance over the Number of RIS elements $M$
Fig. 3 shows the total BS transmit power $P_{\text{tot}}$ vs. the number of RIS elements $M$ by different methods, for the target SINR $\gamma=10$ dB. We see that the BS transmit power for random RIS, direct AO, and IP Method remain roughly flat as $M$ increases. This indicates that they do not provide an effective RIS beamforming solution. In contrast, under Algorithm 1, the resulting BS transmit power is reduced substantially by $4$ dB as $M$ increases from 25 to 300. RIS-SCA has a nearly identical performance to Algorithm 1, while RIS-SDR performs worse than Algorithm 1 with a noticeable gap as large as $2$ dB. This is expected, as SDR is an approximation method with a degraded performance when there are many constraints, and the problem size is relatively large.
We also compare the average computation time of different algorithms in Table I. The computational advantage of Algorithm 1 over RIS-SDR, RIS-SCA, and IP Method as the number of RIS elements $M$ increases is clearly observed, demonstrating the scalability of Algorithm 1 over $M$. In particular, the computational complexity of Algorithm 1 only grows mildly with $M$. Thus, Algorithm 1 provides both the best performance and the fastest computational time than the other three methods, especially for the large value of $M$.
VII-B2 Performance over Target SINR $\gamma$
Fig. 4 shows $P_{\text{tot}}$ vs. target SINR $\gamma$ by different methods. We see that direct AO, Random RIS, and IP method provide either no or only very small improvement over the case of No RIS. A substantial gain can be achieved when the RIS passive beamforming is effectively optimized to improve the channel environment between the BS and users. In particular, our proposed Algorithm 1 provides about $4$ dB power reduction over No RIS for $\gamma\geq 4$ dB. Similar to Fig. 3, RIS-SCA has nearly identical performance to Algorithm 1, while RIS-SDR performs worse than Algorithm 1 with a noticeable $2\sim 3$ dB gap.
Figure 4: Total transmit power $P_{\text{tot}}$ vs. target SINR $\gamma$ ($N=16,M=100$, $(G,K)=(2,2)$).
VII-C Performance Comparison for the MMF Problem
We now evaluate the proposed RIS-assisted multicast beamforming design for the MMF problem. For performance comparison, besides our proposed Algorithm 2, we consider the following methods:

### v3

@@ VI Simulation Results
We consider a downlink RIS-assisted transmission scenario with a BS and an RIS. The locations of the BS and the RIS using the $(x,y,z)$-coordinates in meters are $(x_{\text{BS}},y_{\text{BS}},z_{\text{BS}})=(0,0,20)$ and $(x_{\text{RIS}},y_{\text{RIS}},z_{\text{RIS}})=(70,70,20)$, respectively. The RIS is a rectangular surface with the array elements placed on the $(y,z)$-plane. The users are located randomly within a rectangle area $[45,65]\times[75,125]$ on the $(x,y)$-plane, each at location $(x_{ik},y_{ik},0)$. We set the target SINR for each user the same: $\gamma_{ik}=\gamma$, $\forall i,k$. Unless specified otherwise, our default system setup is given as $N=32$ antennas at the BS, and $M=10\times 10$ reflective elements at the RIS, $G=2$ groups, $K_{i}=K=3$ users, $\forall i$, and SINR target $\gamma=10~{}$dB.
@@
We consider the system operates at carrier frequency $f_{c}=3.5$ GHz. We assume a line-of-sight (LOS) channel between the BS and the RIS and model ${\bf H}_{\text{r}}$ as a Rician fading channel matrix given by ${\bf H}_{\text{r}}=\beta^{\text{BR}}\big{(}\sqrt{\frac{K_{\text{B}}}{1+K_{\text{B}}}}{\bf H}_{\text{r}}^{\text{LOS}}+\sqrt{\frac{1}{1+K_{\text{B}}}}{\bf H}_{\text{r}}^{\text{NLOS}}\big{)}$, where Rician factor $K_{\text{B}}=10$, $\beta^{\text{BR}}$ is the path gain modeled as $\beta^{\text{BR}}\text{[dB]}=-30-22\log 10(d_{\text{B-R}})$ with $d_{\text{B-R}}$ being the BS-RIS distance in meters, ${\bf H}_{\text{r}}^{\text{NLOS}}$ is the non-line-of-sight (NLOS) component modeled as $[{\bf H}_{\text{r}}^{\text{NLOS}}]_{ij}\sim{\cal C}{\cal N}(0,{\bf I})$, and ${\bf H}_{\text{r}}^{\text{LOS}}$ is the LOS component. The LOS component ${\bf H}_{\text{r}}^{\text{LOS}}$ is a function of the BS and RIS locations and is modeled as
${\bf H}_{\text{r}}^{\text{LOS}}={\bf b}_{\text{RIS}}(\phi_{2},\theta_{2}){\bf b}_{\text{BS}}(\phi_{1},\theta_{1})^{H}$ (45)
@@
where $\phi_{1}$ and $\theta_{1}$ are the azimuth and elevation angles of departure (AoD) from the BS to the IRS, and ${\bf b}_{\text{BS}}(\phi_{1},\theta_{1})$ is the BS antenna array 3D steering vector. Assume the BS antenna elements are along the $x$-coordinate, the $n$th entry of ${\bf b}_{\text{BS}}(\phi_{1},\theta_{1})$ is given by
$[{\bf b}_{\text{BS}}(\phi_{1},\theta_{1})]_{n}=e^{-j\frac{2\pi(n-1)\Delta_{\text{BS}}}{\lambda_{c}}\cos(\phi_{1})\sin(\theta_{1})},\ n=1,\ldots,N$
@@
where $\Delta_{\text{BS}}$ is the distance between two adjacent BS antennas, $\lambda_{c}$ is the carrier wavelength, and $\cos(\phi_{1})\sin(\theta_{1})=\frac{x_{\text{RIS}}-x_{\text{BS}}}{d_{\text{B-R}}}$. We assume $\Delta_{\text{BS}}=\frac{\lambda_{c}}{2}$. Similarly, ${\bf b}_{\text{RIS}}(\phi_{2},\theta_{2})$ is the $M\times 1$ RIS element streaming vector, with $\phi_{2}$ and $\theta_{2}$ being the azimuth and elevation angles of arrival from the BS to the RIS, and the $m$th element, for $m\in{\cal M}$, is given by
$[{\bf b}_{\text{RIS}}(\phi_{2},\theta_{2})]_{m}=e^{-j\frac{2\pi\Delta_{\text{RIS}}}{\lambda_{c}}[y_{m}\sin(\phi_{2})\sin(\theta_{2})+z_{m}\cos(\theta_{2})]}$ (46)
@@
where $\Delta_{\text{RIS}}=\frac{\lambda_{c}}{2}$ is the distance between two adjacent RIS elements, $y_{m}=\text{mod}(m-1,M_{y})$ with $M_{y}$ being the number of RIS elements along the $y$-axis, $z_{m}=-\lfloor(m-1)/M_{y}\rfloor$, $\sin(\phi_{2})\sin(\theta_{2})=\frac{y_{\text{BS}}-y_{\text{RIS}}}{d_{\text{B-R}}}$, and $\cos(\theta_{2})=\frac{z_{\text{BS}}-z_{\text{RIS}}}{d_{\text{B-R}}}$.
@@
We assume a LOS channel between the RIS and each user. The channel ${\bf h}^{\text{r}}_{ik}$ between the RIS and user $k$ in group $i$ is generated i.i.d as ${\bf h}^{\text{r}}_{ik}=\beta^{\text{r}}_{ik}\big{(}\sqrt{\frac{K_{\text{r}}}{1+K_{\text{r}}}}{\bf h}_{ik}^{\text{r,LOS}}+\sqrt{\frac{1}{1+K_{\text{r}}}}{\bf h}_{ik}^{\text{r,NLOS}}\big{)}$, where Rician factor $K_{\text{r}}=10$, $\beta^{\text{r}}_{ik}$ is the path gain modeled as $\beta^{\text{r}}_{ik}[\text{dB}]=-30-22\log 10(d^{\text{R-U}}_{ik})$, with $d^{\text{R-U}}_{ik}$ being the distance between the RIS and the user. The NLOS component is ${\bf h}_{ik}^{\text{r,NLOS}}\sim{\cal C}{\cal N}({\bf 0},{\bf I})$. The LOS component ${\bf h}_{ik}^{\text{r,LOS}}={\bf b}_{\text{RIS}}(\phi_{ik}^{\text{r}},\theta^{\text{r}})$ is similar as in (46), where $\phi_{ik}^{\text{r}}$ and $\theta^{\text{r}}$ being the azimuth and elevation AoD from the IRS to each user, with $\sin(\phi_{ik}^{\text{r}})\sin(\theta^{\text{r}})=\frac{y_{ik}-y_{\text{RIS}}}{d^{\text{R-U}}_{ik}}$, and $\cos(\theta^{\text{r}})=\frac{z_{ik}-z_{\text{RIS}}}{d^{\text{R-U}}_{ik}}$.
The channel between BS and user $k$ in group $i$ is modeled as a NLOS channel with Rayleigh fading, which is generated i.i.d as ${\bf h}^{\text{d}}_{ik}\sim{\cal C}{\cal N}({\bf 0},\beta^{\text{d}}_{ik}{\bf I})$, where $\beta^{\text{d}}_{ik}$ is the path gain modeled as $\beta^{\text{d}}_{ik}[\text{dB}]=-32.4-36.7\log 10(d^{\text{B-U}}_{ik})$, with $d^{\text{B-U}}_{ik}$ being the respective BS-user distances. The receiver noise power is set to $\sigma^{2}=-100~{}\text{dBm}$.
@@
VI-A Performance for the QoS Problem
VI-A1 Convergence Behavior
We first study the convergence behavior of the proposed AMBF (Algorithm 1) for ${\cal P}_{o}$. Denote the total transmit power at the BS as $P_{\text{tot}}=\sum_{i=1}^{G}\|{\bf w}_{i}\|^{2}$. Fig. 2(a) shows the trajectory of $P_{\text{tot}}$ over AO iterations by Algorithm 1 for three random channel realizations. We observe that Algorithm 1 converges fast, especially given that fast algorithms are used for each subproblem. Fig. 2(b) shows the convergence behavior of solving subproblem $\widetilde{{\cal S}}_{{\bf w}}$ for RIS phase shift ${\bf e}$ using PSA in the first AO iteration of AMBF. The convergence behavior for three random channel realizations are shown. We see PSA converges after $500\sim 1400$ iterations. Note that the convergence speed becomes faster in the subsequent AO iterations as the algorithm converges. Furthermore, the closed-form updates for PSA in each iteration is computationally inexpensive as shown in the complexity analysis. Thus, the overall computation time is fast despite that the number of iterations can be large especially in the initial AO iteration.
(a) The overall AMBF for ${\cal P}_{o}$.
(b) The subproblem $\widetilde{{\cal S}}_{{\bf w}}$.
Figure 2: Convergence behavior of Algorithm 1 ($N=32,M=100$, $(G,K)=(2,3)$, $\gamma=10$ dB).
VI-A2 Performance Comparison
For performance comparison, we consider the following methods:
@@ VI-A2 Performance Comparison
    DirectSCA [34]: An SCA method to directly solve ${\cal P}_{o}$. Note that this SCA algorithm was developed originally for the RIS-assisted joint beamforming in the multiuser unicast scenario. We can straightforwardly extend it to solve ${\cal P}_{o}$ in the multicast scenario. The SINR constraints w.r.t. ${\bf w}$ and ${\bf e}$ in (2b) are convexified and iteratively solved via a convex solver such as CVX [35]. This method serves as a competitive benchmark.
  4. vi)
    IP Method: Apply the interior-point (IP) algorithm for the joint optimization problem. In particular, we adopt the MATLAB built-in non-linear optimization solver Fmincon to solve ${\cal P}_{o}$.
@@
    DirectAO [29]: An algorithm that directly applies AO to ${\cal P}_{o}$, where the subproblem for ${\bf e}$ is a feasibility problem to satisfy the SINR constraints in (2b) and the unit-modulus constraints in (2c). Each subproblem is solved by an SDR-based algorithm.
Fig. 3 shows the total BS transmit power $P_{\text{tot}}$ vs. the number of RIS elements $M$ by different methods, for the target SINR $\gamma=10$ dB. The performance of random RIS remains roughly flat as $M$ increases, as it does not provide an effective RIS beamforming solution. DirectAO can only provides a small improvement over random RIS, which shows solving the feasibility problem for RIS subproblem does not give an effective solution. Also, the algorithm entails very high computation complexity, and thus its performance is only shown for $M\leq 100$. In contrast, a substantial gain can be achieved when RIS beamforming is effectively optimized. As we see, DirectSCA performs the best with the lowest power. This is expected as it is a joint optimization approach to solve ${\cal P}_{o}$ directly, and SCA performs well in general. However, the main drawback of this approach is the cost of impractically high computational complexity as we will show later. In addition, as $M$ becomes large, the numerical stability of this method is reduced noticeably. In particular, the frequency that the CVX solver fails to find a solution increases to $40\%\sim 50\%$ when $M>100$. In comparison, the performance of the proposed Algorithm 1 closely follows that of DirectSCA, with a gap about $1$ dB. Also, the algorithm always provides a converged solution. IP method for this non-convex problem often has trouble to converge properly, and the resulting performance deteriorates as $M$ increases, with a large gap observed for $M>200$.
Figure 3: Total transmit power $P_{\text{tot}}$ vs. $M$ ($N=32$, $(G,K)=(2,3)$, $\gamma=10$ dB).
@@
$M$ 50 100 200 300 400
Algorithm 1 6.49 17.03 22.2 24.8 37.5
DirectSCA [34] 646 935 2364 2425 4403
IP Method 15.9 19.9 35.1 39.1 47.7
DirectAO 269 1015   –   –    –
We now compare the average computation time of different algorithms in Table I. The computational advantage of Algorithm 1 over DirectSCA, IP Method, and DirectAO is clearly observed as $M$ increases, demonstrating the excellent scalability of Algorithm 1 over $M$. DirectSCA has prohibitively high computational complexity as $M$ increases, making it impractical for real implementation. In contrast, the computational time of Algorithm 1 is several magnitudes lower and only grows mildly with $M$. Thus, Algorithm 1 is a much more efficient algorithm offering competitive performance with significantly lower computational complexity as compared to DirectSCA, especially for the large value of $M$.
Fig. 4 shows $P_{\text{tot}}$ vs. target SINR $\gamma$ by different methods, for $M=256~{}(16\times 16)$. We see that the proposed Algorithm 1 closely follows DirectSCA and provides substantial power reduction over Random RIS and IP method for all values of $\gamma$, with a roughly constant gap.
Figure 4: Total transmit power $P_{\text{tot}}$ vs. target SINR $\gamma$ ($N=32,M=256$, $(G,K)=(2,3)$).
VI-B Performance for the MMF Problem
VI-B1 Convergence Behavior
Fig. 5 shows the convergence behavior of Algorithm 2 for the MMF problem ${\cal S}_{o}$ under three random channel realizations. We see that APSA converges in $1000\sim 2000$ iterations. Since the closed-form updates for APSA in each iteration is computationally inexpensive, the overall computation time is fast despite the number of iterations is large. As we will see later, the overall computational time is low.
Figure 5: Convergence behavior of the Algorithm 2 ($N=32$, $M=100$, $(G,K)=(2,3)$).
VI-B2 Performance Comparison
For performance comparison, besides our proposed Algorithm 2, we consider the following methods: